### v1

# Mutual benefits of social learning and algorithmic mediation for cumulative culture

Mutual benefits of social learning and algorithmic mediation for cumulative culture
Agnieszka Czaplicka1, Fabian Baumann1,2, Iyad Rahwan1
1Center for Humans and Machines,
Max Planck Institute for Human Development, Lentzeallee 94, Berlin 14195, Germany.
2Department of Biology, University of Pennsylvania, Philadelphia, PA 19104
Corresponding author: agaczapl@gmail.com
The remarkable ecological success of humans is often attributed to our ability to develop complex cultural artefacts that enable us to cope with environmental challenges. The evolution of complex culture (cumulative cultural evolution) is usually modeled as a collective process in which individuals invent new artefacts (innovation) and copy information from others (social learning). This classic picture overlooks the fact that in the digital age, intelligent algorithms are increasingly mediating information between humans, with potential consequences for cumulative cultural evolution. Building on an established model of cultural evolution, we investigate the combined effects of network-based social learning and a simplistic version of algorithmic mediation on cultural accumulation. We find that algorithmic mediation has a strong impact on cultural accumulation and that this impact generally increases as social networks become less densely connected. Moreover, cultural accumulation tends to be optimal when social learning and algorithmic mediation are combined, and the optimal ratio depends on the network’s density. Our modeling results are a first step towards formalising the impact of intelligent algorithms on cumulative cultural evolution within an established framework. Models of this kind will also help to uncover mechanisms of human-machine interaction in cultural contexts, guiding hypotheses for future experimental testing.
1 Introduction
Humans’ ability to create complex cultural artefacts such as languages, scientific theories, art and technology, is often seen as one of the most important aspects of our success as a species [1, 2]. An important driving force behind this cultural accumulation is the ability of humans to learn socially, i.e. from their peers [3, 4], and pass our their acquired knowledge on to others [3, 5, 1]. Social learning is a highly adaptive process, influenced by cognitive biases that can manifest as different social learning strategies ranging from copying successful individuals (prestige bias) [6, 7], to frequent cultural artefacts (context bias) [8, 9] to specific information (content bias) [10]. At the population level, social learning is a necessary condition for cumulative cultural evolution, a process that describes the emergence of increasingly complex cultural artifacts through a ratchet effect: By preserving important information over time—i.e., from generation to generation—social learning allows collectives to produce cultural artefacts that lie outside of the scope of individuals [11, 12]. Yet the accumulation of culture in populations can also be limited by factors such as population size [12] or the structure of the underlying social network [13].
In the digital age, much of human social learning is shaped by intelligent algorithms. Instead of learning directly from their peers, individuals absorb information mediated by algorithms. Examples of such algorithms include recommendation systems that influence consumer behavior through personalized product information [14], algorithmically curated content feeds on social media platforms [15] and search engines that rank websites and online content [16]. By facilitating the targeted retrieval of information, these systems can significantly improve the ability of individuals to navigate complex online information landscapes and promote the transfer of knowledge within collectives. However, information-mediating algorithms also have a potential downside: algorithmic personalization has been shown to restrict the scope for individual exploration and limit the diversity of information, for example by actively creating filter bubbles [17] or exacerbating natural tendencies towards homophily [18, 19]. Recognizing these impacts, previous theoretical and empirical works have investigated the impact of algorithms on the diversity of content [20], the formation of opinions [21, 22], and the spread of misinformation [23]. However, little attention has been paid to the question of how algorithmic mediation may affect fundamental and more general processes of cultural evolution [24, 25]. In particular, it is unclear how algorithmic mediation interacts with human social learning to influence the accumulation of complex cultural artifacts.
In this paper, we take a step forward to close this gap by building on an established model of cultural evolution [26]. More specifically, we extend the model by a simple algorithm that first aggregates all available information and then redistributes it in a personalized way. This allows us to study the interplay between social learning (i.e., people share information based on a social network) and algorithmic mediation (i.e., information is mediated by an algorithm). Our modeling results shed light on an emerging trade-off that humans face in the digital age: cultural accumulation is neither optimal through pure social learning nor through purely algorithmically mediated information transfer. Instead, both types of learning are most powerful when combined, where the optimal ratio depends on the underlying social network. The dynamic interplay between social learning and algorithmic mediation suggests that understanding cultural evolution in the digital age requires considering both types of learning together, rather than in isolation.
2 Model
We build on previous theoretical work [26] that examined how cultural accumulation is affected by different constraints on the individual level. In particular, while the impact of different (social) learning strategies and the costs associated to learning on cultural accumulation were studied, the influence of a social network structure and algorithmic mediation was not investigated.
To fill this gap, we first consider cultural accumulation in populations of agents that communicate based on random (social) networks with different densities of network connections. In a second step, we extend the model by a simple algorithm that mediates information among agents. The individual-based model presented in [26] considers a dynamic process that evolves in generations, with each ”time step” comprising the birth and death of an entire generation of agents. First, the agents copy from the previous generation and then, having learnt all possible information, they start to innovate. Here we deviate from this approach. We modify the model by re-scaling the dynamics so that at each time step a single cultural trait can be acquired either by copying or by innovation [27]. This modification allows us to study the interplay between different mechanisms of information transmission between agents—social learning and algorithmic mediation—and the consequences for population-scale cultural accumulation.
2.1 Cultural space
We consider a space of cultural traits (or cultural space) that defines how new traits (innovations) arise based on existing ones. In line with previous works, the cultural space consists of several cultural branches (or paths), and that traits are functionally dependent, reflecting the cumulative nature of human culture [13, 28, 26, 29].
More specifically, the cultural space is structured as follows. Each trait $x_{l}$ is defined by two components: the cultural branch $x$ it belongs to and its complexity level $l$, see Fig. 1A. Cultural branches $x$ represent the fact that cultural evolution often follows multiple parallel paths of development [28, 13], and levels $l$ model the functional dependency and increasing complexity of cultural artifacts built on top of each other. We consider a finite set of possible branches, i.e., $x=1,...,X$ and assume that cultural complexity levels $l$ are unbounded, i.e. $l=1,2,...,\infty$. The functional dependency of the traits $x$ additionally requires that the traits are acquired in ascending order, starting with the lowest level $l=1$, or more generally: an agent that wants to acquire a trait on level $l$ must possess a trait on level ${l-1}$.
Each cultural trait $x_{l}$ is characterised by an inherent quality score, or payoff, $z(x_{l})$, and we assume that the extent to which the traits are beneficial to agents increases with $z$ [5, 13, 30, 31]. Note, however, that we do not link an individual’s probability to survive to their payoff. Instead, cultural selection happens through selective social learning, where agents copy from agents with high payoffs [30, 32]. Motivated by empirical research on scientific [33] and technological [34] breakthroughs and previous studies on cultural evolution [5], we assume that high-quality (i.e. high-payoff) innovations are rare among many low-quality attempts to innovate. Hence, we draw the payoffs of newly innovated items, $z(x)$, from an exponential distribution, and subsequently square, double, and finally round values to an integer value [27]. This procedure results in a few traits with a payoff of around $z=50$, and roughly half of the traits with $z=0$ [26].
2.2 Dynamics
We consider a population of $N$ agents that are embedded in a social network. Initially, agents are naive, i.e. they do not possess any cultural traits, and their individual effort budget is $B$. A finite effort budget $B$ models the fact that individuals have limited resources, which also constrains cumulative cultural evolution at the collective level. In particular, over time agents spend their budget $B$ on actions to acquire cultural traits, either by copying already discovered traits (learning), or by discovering new traits (innovation). The dynamics of a focal agent $i$ at time $t$ (at development level $l_{i}$ and remaining resources $B(t)$) can be described as a cycle of the following three events:
$(i)$ Death-birth process.
With probability $q=1/N$ agent $i$ is replaced by a naive agent, $i^{\prime}$, which corresponds to the replacement of on average one agent per time step. The initial effort budget of agent $i^{\prime}$ is $B$ and they do not possess any cultural traits. Agent $i^{\prime}$ takes the network position of agent $i$.
$(ii)$ Social learning.
Agent $i$ selects its neighbour $j$ with the highest cumulative payoff [7]. If agent $j$ is at a higher development level than agent $i$, i.e. $l_{j}>l_{i}$, agent $i$ copies the $(l_{i}+1)$-stage trait of agent $j$, and the learning cost $C_{s}$ is subtracted from agent $i$’s budget. If the copying of a trait is not successful, either because the resources of agent $i$ are not sufficient ($B(t)<C_{s}$) or because the level of the highest trait of agent $j$ is too low, i.e. $l_{j}<l_{i+1}$, the learning costs $C_{s}$ are not deducted from the agent’s budget. Agent $i$ then continues with the innovation step (see below). If the social learning step was successful, the innovation step is skipped. Social learning is schematically depicted in Fig. 1C.
$(iii)$ Innovation.
If social learning was unsuccessful, and agent $i$ has a sufficient budget ($B(t)\geq C_{i}$), it will proceed with innovation. First, $i$ randomly selects a cultural branch $x$ and attempts to innovate a trait on level $l_{i}+1$. If the trait is viable, i.e. its payoff $z(x_{l_{i}+1})>0$, the innovation step is successful and agent $i$ acquires the trait. Note that unsuccessful innovation concludes the time step for agent $i$ without acquiring a new item. Regardless of the success of the innovation step, the cost of innovation $C_{i}$ is subtracted from agent $i$’s effort budget. Innovation is schematically depicted in Fig. 1B.
To quantify the performance of a population with respect to cultural accumulation, we calculate the mean payoff of all agents in the population, i.e. $\bar{Z}(t)=N^{-1}\sum Z_{i}(t)$, where $Z_{i}$ is the cumulative payoff of agent $i$, defined as $Z_{i}=\sum_{l=1}^{l_{\mathrm{max}}}z(x_{l})$. Note that to compute $Z_{i}$, we only sum over the complexity level $l$, as a consequence of the assumed dynamics of innovation and learning: an agent can only acquire a maximum of one trait per complexity level $l$. Additionally, we assume that the cost of acquiring a single trait depends only on the action performed, i.e., learning or innovation, but is independent of other factors, including the payoff $z$ and the complexity level $l$ of the trait. Furthermore, we follow the intuition that it is easier to copy something from a peer than to innovate it, i.e. we assume $C_{s}<C_{i}$ [26].
2.3 Algorithmic mediation
To investigate how algorithmic mediation can affect cultural accumulation, and to shed light on the interplay between different types of information propagation, i.e. social and algorithm-mediated learning, we extend the model described in the previous section. While still being embedded in a social network, agents can now also learn cultural traits through algorithmic mediation. For simplicity, we implement algorithmic mediation as a ’perfect’ recommender system that has full knowledge about all cultural traits innovated so far in the system, i.e. the cultural reservoirs of every agent. The knowledge about traits’ development levels $(l)$ and their payoffs ($z(x_{l})$) allows the algorithm to provide individuals with personalized recommendations. We assume that learning through algorithmic mediation requires less effort than social learning, i.e., we assume $C_{r}<C_{s}$, where $C_{r}$ defines the costs of algorithmic mediation. In each time step, the agent copies a trait through algorithmic mediation with probability $r$, or otherwise, with probability ($1-r$), the agent learns socially.
Figure 1: Schematic depiction of key model components. A: Open-ended cultural space depicted with four branches ($X=4$), and an infinite number of (complexity) levels $l$. B: Agents innovate a new item on a randomly chosen branch on a level $l+1$. C: Social learning is modeled as a two-step process. (1) Agents choose their neighbor $j$ with the highest cumulative payoff $Z_{j}$. (2) If agent $j$ possesses an item on a sufficient level ($l+1$), the item is copied by agent $i$ and the cost $C_{s}$ is paid. D: Algorithmic mediation is also modeled as a two-step process. (1) The recommender system aggregates all existing cultural traits and sorts traits on level ($l+1$) according to their payoffs $z$. (2) Agent $i$ then receives a recommendation of the trait with the highest payoff.
Algorithmic recommendation (with probability $r$).
In every time step, the recommender system aggregates all cultural traits and selects only those on level $l_{i}+1$ (personalization). After sorting those $l_{i}+1$-level traits in descending order (according to their payoffs) it recommends the highest-payoff trait, which is acquired by agent $i$ and the cost $C_{r}$ is deducted from its budget. Note that when no suitable traits can be recommended, or if agent $i$ does not have enough resources ($B(t)<C_{r}$), the recommendation fails, the agent does not pay a cost, and continues with the innovation step. When the recommendation is successful, agent $i$ does not attempt to innovate. Algorithmic mediation is schematically depicted in Fig. 1D.
Social learning (with probability $1-r$).
Agent $i$ attempts to copy a cultural trait from neighbors (the social learning step is identical to the one described in the previous section).
The order of events in each time step is identical to the one described in the previous section, i.e., consisting of $(i)$ the death-birth process, $(ii)$ learning (either socially or through algorithmic mediation), and $(iii)$ innovation. The different model components are schematically depicted in Fig. 1.
3 Results
If not indicated otherwise, we present results for populations of $N=100$ agents that are embedded in a random social network, where each pair of individuals is connected with probability $p$ and thus has on average $k=pN$ neighbors. The effort budget of agents is set to $B=1000$, the costs for acquiring cultural traits are fixed at $C_{i}=10$, $C_{s}=5$, and $C_{r}=1$, and we assume $X=100$ cultural branches.
Due to the finite effort budget $B$ and the low probability of the death-birth process $q$, the system reaches a quasi-stationary state regardless of the other parameters. In the following, we will thus focus on these ”maximally achievable values”, i.e., population mean values at the stationary state. For brevity, we will refer to these values as ”mean group values” and mark them with a subscript max. Each simulation continues for $t=20N$ time steps and the last $200$ time steps are used to compute the mean group values. Our results are then reported as averages over $200$ realizations.
3.1 Network-based social learning
Figure 2 depicts the results for fully connected networks ($p=1$). Similar to the findings reported in [26], we observe that after the initial growth of $\bar{Z}$ the dynamics reach a quasi-stationary state where the mean group payoff approaches its maximum value, denoted as $\bar{Z}_{\mathrm{max}}$. In line with classical results in cultural evolution [12], we find that the maximum cultural complexity increases with the system size uncovering a logarithmic relationship. This relationship becomes more clear by plotting $\bar{Z}_{\mathrm{max}}$ as a function of $\log(N)$ (system size), as depicted in the right panel of Figure 2.
Figure 2: (left panel) The mean group payoff for fully connected networks as a function of time for different system sizes. (right panel) Maximally achievable group payoff as a function of system size. The remaining model parameters are: $B=1000$, $C_{i}=10$, $C_{s}=5$, $X=100$.
Figure 3 shows the results for random networks with different connectivity, as parameterized by the average number of neighbours $k$ of an agent. From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of cultural traits $\bar{T}_{\mathrm{max}}$, mean group payoff $\bar{Z}_{\mathrm{max}}$. The mean group level grows with the number of neighbors until it reaches a plateau for very dense networks. Instead, the number of discovered items $\bar{T}_{\mathrm{max}}$ decreases rapidly and monotonically with $k$. The most striking behavior is observed for the mean group payoff. Initially, for very sparse networks (low values of $k$) we observe an increase of $\bar{Z}_{\mathrm{max}}$with $k$, but after a maximum value is reached, for intermediate levels of connectivity, we observe that $\bar{Z}_{\mathrm{max}}$ decreases with $k$.
Figure 3: System behavior in a stationary state as a function of network connectivity $k$. From left to right we plot mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$. The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $X=100$
3.2 Algorithmic mediation
Next, we focus on the quantitative and qualitative impact of the algorithmic mediation on cumulative cultural evolution. The results are shown in Fig. 4. Only the mean group level, $\bar{L}_{\mathrm{max}}$, shows a monotonous increase with increasing $r$ for every value of $k$, see left panel of Fig. 4. In line with the observations from the previous section, a higher level of development is observed for denser networks. By contrast, the mean group payoff $\bar{Z}_{\mathrm{max}}$ gives rise to maximum values for intermediate levels of $r$, and these maximum values decrease with $k$. This suggests a trade-off between network connectivity and the ratio of recommendations, which can be seen more clearly in Fig. 5. The highest payoff is observed for very sparse networks (every agent has only one neighbor on average) and relatively high levels of recommendations ($r\approx 0.7$), which might be due to the high number of innovated traits $\bar{T}_{\mathrm{max}}$, for low values of $k$, as can be seen from the middle panel of Fig.  4). With fewer neighbors an individual benefits from a higher number of recommendations, but this number decreases with $k$. Note that both limiting cases of $r=0$ and $r=1$ never give rise to maximum values, independently of the connectivity, suggesting that network-based social learning and algorithmic-mediation of information are mutually beneficial.
Figure 4: Stationary states for mean group level $\bar{L}_{\mathrm{max}}$, mean number of traits $\bar{T}_{\mathrm{max}}$, and mean group payoff $\bar{Z}_{\mathrm{max}}$ (from left to right) as a function of recommendation probability $r$, and for different network connectivities $k$ including complete graphs (CG). The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
Figure 5: Heatmap of the mean group payoff $\bar{Z}_{\mathrm{max}}$ as a function of the average connectivity $k$ and probability of reccomendation $r$. The remaining model parameters are: $N=100$, $B=1000$, $C_{i}=10$, $C_{s}=5$, $C_{r}=1$, $X=100$
From our numerical simulations, it appears that both the size of the system ($N$) and the number of possible development paths $X$, do not qualitatively affect our modeling results. We also did not observe non-trivial relationships between the parameters related to learning costs (i.e. $C_{i}$, $C_{s}$, $C_{r}$, $B$) and the measures of cultural accumulation. Intuitively, the more we increase the cost of learning or innovation, the lower the payoff of the group. Similarly, a larger effort budget translates into higher cultural accumulation, where the resulting dynamics depend on the network topology and the recommendation probability $r$. Nevertheless, our findings are applicable within the context of reasonable parameter values, where the costs of acquiring items are significantly lower than the effort budget ($C\ll B$). Finally, it is important to point out that although we focused on cases in which the costs of social learning are higher than the costs of algorithmic mediation, we confirmed that our results are robust for $C_{s}=C_{r}$.
4 Discussion
We investigated the effects of network-based social learning and algorithmic mediation on cultural accumulation by building on, and extending, previous modeling efforts from the field of cultural evolution [26]. In the first part, we considered the model without algorithmic mediation and found a strong dependence of cultural accumulation on population size. In particular, for fully connected networks, the ability of a population to accumulate culture showed a positive relation (logarithmic growth) with population size, which is consistent with established findings in cultural evolution [6, 35, 36, 37, 38, 39]. We then investigated cultural accumulation in random networks that differ in the density of connections, as quantified by the average number of neighbors $k$ of a single agent. Although increasing the link density always led to higher levels of development ($\bar{L}_{\mathrm{max}}$), this was not the case for the mean group payoff ($\bar{Z}_{\mathrm{max}}$) and the number of innovated traits ($\bar{T}_{\mathrm{max}}$), which peak at intermediate values of $k$ and monotonously decrease with $k$, respectively. These non-trivial behaviors indicate that cultural accumulation is subject to an exploration-exploitation trade-off between innovation (exploration) and social learning (exploitation), a phenomenon that was previously also observed in both experimental and theoretical works on social dynamics including cultural evolution [13, 28], wisdom of the crowds [40], as well as collective problem-solving [41, 31, 42, 43]. Dense random networks (i.e. networks with high values of $k$) generally allow to disseminate information about cultural traits efficiently (exploitation) and therefore lead to high cultural complexity, however, do so at the expense of cultural diversity (number of traits) as the rate of innovation is impaired due to an over-reliance on social learning. Furthermore, previous empirical and theoretical findings in cultural evolution and social learning studies suggest that collective performance depends on both the complexity of the task and the underlying network structure. In particular, it was found that performance in dense networks is often decreased for complex tasks [31, 42, 43], a phenomenon which is best explained by sufficient levels of transient diversity in less densely connected systems [44]. In Fig. 3, the shapes of $\bar{L}_{\mathrm{max}}$ and $\bar{Z}_{\mathrm{max}}$, as a function of $k$, therefore indicate that the different sub-processes of cultural accumulation vary in complexity: While reaching a high level of complexity is a rather simple problem (high value of $\bar{L}_{\mathrm{max}}$ in dense networks), it is more complex to collectively coordinate on high average payoffs (low values of $\bar{Z}_{\mathrm{max}}$ in dense networks). In other words, populations easily reach high level of development, but do not necessarily do so with high collective payoffs. In our model, collective payoffs are optimal at a medium level of connectivity, where agents optimally solve the explore-exploit trade-off, i.e., have access to the right amount of information through social learning while maintaining sufficient incentives to innovate.
In the second part, we investigated the combined impact of network-based social learning and algorithmic mediation on cultural accumulation. For a fixed value of $k$, we find that increasing $r$—the ratio between algorithmic mediation over network-based social learning—strongly affects the overall dynamics: while the developmental level monotonously increases with $r$, the mean payoff gives rise to maximum values for intermediate levels of algorithmic mediation, which suggests a non-trivial interaction between $k$ and $r$. The less we are connected, the more we can benefit from algorithmic mediation. For a given link density, however, the benefits of increased information sharing via algorithmic mediation become outweighed by their detrimental consequences: too high levels of $r$ create too much of effective connectivity which decreases performance as $r\rightarrow 1$. This is in line, with the strongly decreasing number of traits as $r$ increases, indicating low levels of innovation. These results make a clear case for studying the dynamics of algorithmic mediation not in isolation but in relation to social learning strategies motivated by cognitive science (e.g. payoff-biased learning). In our model, both algorithmic mediation and social learning acting alone consistently lead to worse payoffs and lesser cultural traits than a combination of both processes. This behavior may point toward a general trade-off in socio-technical systems between access to information (and thus the possibility to exchange it) and the ability of individuals to innovate: while algorithms can be beneficial for cultural accumulation, they may also lead to detrimental effects if used excessively [45, 46].
Our research takes a first step in modeling the impact of personalized algorithmic mediation on cumulative cultural evolution, however, it does not come without limitations. By focusing on the algorithmic mediation of information, inspired by recommender systems or feed algorithms, we have not considered other types of intelligent algorithms that are increasingly intertwined with human culture, in particular, generative artificial intelligence [47]. Importantly, generative AI not only re-distributes information, but is able to create novel cultural variation, which is not captured in our models. Furthermore, in line with previous models, we considered a very simplistic cultural space, namely cultural branches that are independent from each other. A more realistic approach could therefore take into account that cultural innovations often arise through the recombination of artefacts from distant domains, i.e. modeling interdependencies between branches. Similarly, by investigating random networks and agents with identical social learning strategies we have omitted many of the behavioral and structural heterogeneities observed in real social systems [48, 49]. The use of more complex network models—such as scale-free [50] and small-world networks [51]—as well as agents with individual preferences are promising avenues for future research directions. Finally, we have operationalised algorithmic mediation as a simple system that provides agents with ”ideal” personalised recommendations. Therefore, extending our approach to more realistic models of algorithmic mediation, e.g. recommendations based on collaborative filtering [52], is a promising next step. Overall, our work is only a first step towards a better theoretical understanding of the impact of algorithms on cumulative cultural evolution. Ultimately, the predictions of our model must stand up to empirical testing, both through controlled laboratory experiments and real-world observational studies across different online environments.
Code availability
Simulation code is available at https://github.com/aga-cz/Mutual-benefits-of-SL-and-RS